OpenStudy (anonymous):

Find the linearization of f(x) = x^3 + x at a = 2, and then use it to estimate f(2.2).

OpenStudy (anonymous):

take derivative of x^3+x

OpenStudy (anonymous):

3x^2+1?

OpenStudy (anonymous):

good. L(x)=f(x)+f'(x)dx f(x)=x^3+x f'(x)=3x^2+1

OpenStudy (anonymous):

evaluate at 2

OpenStudy (anonymous):

37?

OpenStudy (anonymous):

I think that's wrong :/ Ughh. Im so dumb

OpenStudy (anonymous):

f'(2)=4 f(2)=2^3+2=10 dx=0.2 because we want to estimate 0.2( 2.2-2) f(2.2)=10+4(0.2)

OpenStudy (anonymous):

then we add them?

OpenStudy (anonymous):

10+4(0.2)=10.8

OpenStudy (anonymous):

Its telling me the answer has to be between 12 and 14 :?

OpenStudy (anonymous):

f'(2)=13 (arithmethic error above) f(2)=2^3+13(.2)=10 dx=0.2 because we want to estimate 0.2( 2.2-2) f(2.2)=10+13(0.2)=12.6

OpenStudy (anonymous):

what went wrong though? I cant find the error..

OpenStudy (anonymous):

and thanks :) that answer is correct but what was wrong in our old answer?

OpenStudy (anonymous):

plugging in 2 into 3x^2+1, I got 4 instead of correct answer of 13; silly arithmetic error